# The gauge coupling evolutions of an ${\rm SU}(8)$ theory with the maximally symmetry breaking pattern

### v1

@@ 1 Introduction
Grand Unified Theories (GUTs), with their original formulations based on the gauge groups of $\rm SU(5)$ [1] and $\rm SO(10)$ [2], were proposed to not only unify all three fundamental symmetries into one simple Lie group of $\mathcal{G}_{U}$, but to unify all Standard Model (SM) fermions into some anomaly-free irreps of the $\mathcal{G}_{U}$. The well-known minimal setup include the Georgi-Glashow $\rm SU(5)$ GUT [1] with $3\times\left[\mathbf{\overline{5_{F}}}\oplus\mathbf{10_{F}}\right]$, and the Fritzsch-Minkowski $\rm SO(10)$ GUT [2] with $3\times\mathbf{16_{F}}$. Meanwhile, these $\rm SU(5)$ and/or $\rm SO(10)$ GUTs, with or without their supersymmetric (SUSY) extensions, are far from satisfactory, given that three-generational SM fermion masses and the Cabibbo-Kobayashi-Maskawa (CKM) mixing pattern [3, 4] were not ultimately explained.
@@
In a pioneering work by Georgi [5], an extension to the gauge group of ${\rm SU}(N>5)$ was proposed to embed three-generational SM fermions non-trivially. In his original proposal, any anti-symmetric rank-$k$ fermion irrep can appear at most once with the anomaly-free condition. Physically, this leads to a non-repetitive structure for both left-handed and right-handed components of chiral fermions based on an ${\rm SU}(11)$ gauge group. Given the current LHC measurements of one single SM Higgs boson so far, which have already confirmed the hierarchical Yukawa couplings of the third-generational SM quarks/leptons and displayed the evidence for the muon [6, 7], Georgi’s original proposal is quite insightful and can be further relaxed by requiring distinctive symmetry properties of three generations in the UV theory. A minimal ${\rm SU}(8)$ theory [8, 9, 10] was suggested to embed three-generational SM quarks/leptons non-trivially  [^1] By decomposing the ${\rm SU}(8)$ fermionic irreps into the ${\rm SU}(5)$ irreps, one finds three identical SM $\mathbf{\overline{5_{F}}}$’s and three distinctive $\mathbf{10_{F}}$’s. Hence, it is sufficient to obtain three distinctive SM generations in the UV.. The ${\rm SU}(8)$ symmetry breaking pattern includes three intermediate scales above the EW scale. By analyzing the emergent non-anomalous global $\widetilde{{\rm U}}(1)_{T}$ symmetry in the UV, we derived the non-anomalous global $B-L$ symmetry in the SM, and one unique SM Higgs doublet from the $\mathbf{70_{H}}$ of the minimal setup was conjectured based on the requirement of the vanishing global $B-L$ charge [9]. This SM Higgs doublet was found to give natural top quark mass with the renormalizable Yukawa coupling at the tree level. Due to the distinctive symmetry properties of three-generational SM fermions in the UV, we found in Ref. [10] that all SM quark/lepton masses as well as the CKM mixing pattern can be explained with (i) both $d=5$ direct Yukawa couplings and indirect Yukawa couplings to the $d=5$ irreducible Higgs mixing operators induced by the inevitable gravitational effects, (ii) three reasonable intermediate symmetry-breaking scales, and (iii) precise identifications of the SM flavors in their UV irreps.
@@
As a unified gauge theory, a central issue is to achieve the gauge coupling unification in terms of their renormalization group equations (RGEs)  [^2] Some early studies of the RGEs based on the ${\rm SU}(5)$ and ${\rm SO}(10)$ groups include Refs. [11, 12, 13].. A strong-weak-weak (SWW) symmetry breaking pattern of
@@
were predicted with the reduced Planck scale of $M_{\rm pl}=(8\pi G_{N})^{-1/2}=2.4\times 10^{18}\,{\rm GeV}$. The maximally broken subgroup of $\mathcal{G}_{441}$ further admits the $\mathcal{G}_{\rm SM}$ as its subgroup  [^3] Other different symmetry breaking pattern of the ${\rm SU}(8)$ group was previously described in Ref. [14].. These scales were obtained by the reasonable predictions of three-generational SM quark/lepton masses and the CKM mixing pattern, together with the precise identifications of the SM flavors in their UV irreps. In this paper, we perform the RG analysis of the gauge couplings in the extended strong and weak sectors based on the benchmark point in Eq. (2).
The rest of the paper is organized as follows. We review the ${\rm SU}(8)$ theory in Sec. 2, with the focus on the symmetry breaking pattern. In Sec. 3, we define the gauge sector of the ${\rm SU}(8)$ theory along its symmetry breaking pattern defined in Sec. 2. The gauge couplings at each symmetry breaking stage will be defined. Due to the intrinsic symmetry properties, we explicitly display the flavor non-universalities of the $\mathcal{G}_{441}$ and $\mathcal{G}_{341}$ theories as well as the flavor universality of the $\mathcal{G}_{331}$ theory through their neutral currents. In Sec. 4, we study gauge coupling unification in terms of the two-loop RGEs. The ${\rm SU}(8)$ theory with the minimal fermion/Higgs contents as well as the intermediate symmetry breaking scales in Refs. [9, 10] cannot achieve the gauge coupling unification in terms of the conventional field theoretical analyses. Instead, inclusion of additional ${\rm SU}(8)$ adjoint Higgs fields and the $d=5$ Hill-Shafi-Wetterich (HSW) operator [15, 16] are suggested to achieve the gauge coupling unification within the field theoretical framework. We summarize and make future perspective in Sec. 5, such as the possible issues in the supersymmetric extensions. An App. A is given to define the gauge group indices and the decomposition rules according to the symmetry breaking pattern in the ${\rm SU}(8)$ theory. App. B is added to prove a unique feature of the ${\rm U}(1)$ charge definition in the maximally symmetry breaking pattern in Eq. (1).
2 The ${\rm SU}(8)$ theory and its symmetry breaking pattern
@@ 2.1 The setup of the ${\rm SU}(8)$ theory
The ${\rm SU}(8)$ theory contains two following chiral IRAFFSs at the GUT scale
@@
with undotted/dotted indices for the $\mathbf{\overline{8_{F}}}$’s in the rank-$2$ IRAFFS and the rank-$3$ IRAFFS, respectively. The Roman numbers and the Arabic numbers are used for the heavy partner fermions and the SM fermions, respectively. These three $\mathbf{10_{F}}$’s are transforming differently, while three identical $\mathbf{\overline{5_{F}}}$’s can be identified from nine $\mathbf{\overline{8_{F}}}$’s, which are sufficient to guarantee that three-generational SM quarks and leptons must transform differently in the UV theory.
The non-anomalous global DRS symmetries from fermions in Eq. (3) are
@@
and we also denote the anomalous global Peccei-Quinn symmetries [17] as
@@
Both $\widetilde{{\rm U}}(1)_{T_{2}}$ and $\widetilde{{\rm U}}(1)_{{\rm PQ}_{2}}$ can be thought as the linear combinations of two global $\widetilde{{\rm U}}(1)$’s of the $\mathbf{\overline{8_{F}}}^{\omega}$ and the $\mathbf{28_{F}}$ in the rank-2 IRAFFS, and similarly for the $\widetilde{{\rm U}}(1)_{T_{3}}$ and $\widetilde{{\rm U}}(1)_{{\rm PQ}_{3}}$ symmetries in the rank-3 IRAFFS. Since the global $B-L$ symmetry should be identical for all three generations, we further require a common non-anomalous $\widetilde{{\rm U}}(1)_{T}\equiv\widetilde{{\rm U}}(1)_{T_{2}}=\widetilde{{\rm U}}(1)_{T_{3}}$ between two IRAFFSs. Likewise, we also assume a common anomalous $\widetilde{{\rm U}}(1)_{\rm PQ}\equiv\widetilde{{\rm U}}(1)_{{\rm PQ}_{2}}=\widetilde{{\rm U}}(1)_{{\rm PQ}_{3}}$ between two IRAFFSs in Eq. (5). Accordingly, the non-anomalous global $\widetilde{{\rm U}}(1)_{T}$ charges and the anomalous global $\widetilde{{\rm U}}(1)_{\rm PQ}$ charges for fermions are assigned in Tab. 1.
@@
The $\mathbf{63_{H}}$ can be decomposed under the $\mathcal{G}_{441}$ as follows
@@
where the VEV is developed by the $(\mathbf{1}\,,\mathbf{1}\,,0)_{\mathbf{H}}$ component, and the $(\mathbf{\overline{4}}\,,\mathbf{4}\,,+\frac{1}{2})_{\mathbf{H}}\oplus(\mathbf{4}\,,\mathbf{\overline{4}}\,,-\frac{1}{2})_{\mathbf{H}}$ represent the Nambu-Goldstone bosons. We assign the non-anomalous global $\widetilde{{\rm U}}(1)_{T}$ charges and the anomalous global $\widetilde{{\rm U}}(1)_{\rm PQ}$ charges for Higgs fields in Tab. 1 according to the renormalizable Yukawa couplings in Eq. (6). The anomalous global $\widetilde{{\rm U}}(1)_{\rm PQ}$ charges are assigned such that
$p:q_{2}\neq-3:+2\,,\quad p:q_{3}\neq-3:+1\,.$ (10)
Obviously, the only non-renormalizable term in Eq. (6) also has a non-vanishing global $\widetilde{{\rm U}}(1)_{\rm PQ}$ charge according to Tab. 1 and Eq. (10). Thus, it can be explicitly broken by the gravitational effect.
@@ 2.2 The SWW symmetry breaking pattern of the ${\rm SU}(8)$ theory
According to Ref. [18], only the adjoint Higgs field can achieve the maximal symmetry breaking pattern of ${\rm SU}(N)\to{\rm SU}(N-k)\otimes{\rm SU}(k)\otimes{\rm U}(1)$ (with $k=[\frac{N}{2}]$). The SWW symmetry breaking pattern in Eq. (1) follows from Ref. [10] will be further studied in the following discussions. The GUT scale symmetry breaking will be due to the VEV of the ${\rm SU}(8)$ adjoint Higgs field. In the subgroup of $\mathcal{G}_{441}$, both the strong and the weak symmetries are extended beyond the SM gauge symmetries. All intermediate symmetry breaking stages will be due to the VEVs of the Higgs fields from the gauge-invariant Yukawa couplings of the ${\rm SU}(8)$ theory. After the first stage of symmetry breaking pattern of $\mathcal{G}_{441}\to\mathcal{G}_{341}$ in Eq. (1), the sequential symmetry breaking stages are uniquely determined. Along this symmetry breaking pattern, the non-anomalous global $\widetilde{{\rm U}}(1)_{T}$ symmetry will evolve to the global $\widetilde{{\rm U}}(1)_{B-L}$ at the EW scale according to the following sequence [9]
@@ 2.3 Decompositions of the ${\rm SU}(8)$ fermions
Table 2: The $\rm SU(8)$ fermion representation of $\mathbf{\overline{8_{F}}}^{\Omega}$ under the $\mathcal{G}_{441}\,,\mathcal{G}_{341}\,,\mathcal{G}_{331}\,,\mathcal{G}_{\rm SM}$ subgroups for the three-generational ${\rm SU}(8)$ model, with $\Omega\equiv(\omega\,,\dot{\omega})$. Here, we denote $\underline{{\mathcal{D}_{R}^{\Omega}}^{c}={d_{R}^{\Omega}}^{c}}$ for the SM right-handed down-type quarks, and ${\mathcal{D}_{R}^{\Omega}}^{c}={\mathfrak{D}_{R}^{\Omega}}^{c}$ for the right-handed down-type heavy partner quarks. Similarly, we denote $\underline{\mathcal{L}_{L}^{\Omega}=(\ell_{L}^{\Omega}\,,-\nu_{L}^{\Omega})^{T}}$ for the left-handed SM lepton doublets, and $\mathcal{L}_{L}^{\Omega}=(\mathfrak{e}_{L}^{\Omega}\,,-\mathfrak{n}_{L}^{\Omega})^{T}$ for the left-handed heavy lepton doublets. All left-handed neutrinos of $\check{\mathcal{N}}_{L}$ are sterile neutrinos, which are $\mathcal{G}_{\rm SM}$-singlets and do not couple to the EW gauge bosons.
@@
Table 3: The $\rm SU(8)$ fermion representation of $\mathbf{28_{F}}$ under the $\mathcal{G}_{441}\,,\mathcal{G}_{341}\,,\mathcal{G}_{331}\,,\mathcal{G}_{\rm SM}$ subgroups for the three-generational ${\rm SU}(8)$ model. All SM fermions are marked with underlines.
@@
Table 4: The $\rm SU(8)$ fermion representation of $\mathbf{56_{F}}$ under the $\mathcal{G}_{441}\,,\mathcal{G}_{341}\,,\mathcal{G}_{331}\,,\mathcal{G}_{\rm SM}$ subgroups for the three-generational ${\rm SU}(8)$ model. All SM fermions are marked with underlines.
By following the symmetry breaking pattern in Eq. (1), we tabulate the fermion representations at various stages of the ${\rm SU}(8)$ model in Tabs. 2, 3, and 4. All ${\rm U}(1)$ charges at different stages are obtained according to Eqs. (222a), (222b), (222c), and (222d). According to the rule by Georgi [5], it is straightforward to find $n_{g}=3$ in the current setup. All chiral fermions are named by their $\mathcal{G}_{\rm SM}$ irreps. For the right-handed down-type quarks of ${\mathcal{D}_{R}^{\Omega}}^{c}$, they are named as follows
@@
The flavor indices in Eqs. (2.3) and (2.3) are chosen in accordance with the symmetry breaking pattern depicted in Sec. 2. The SM flavor names in Tabs. 2, 3, and 4 have been determined according to the consistent mass hierarchies and the CKM mixing pattern predicted in Ref. [10].
Through Tabs. 2, 3, and 4, it is straightforward to find three-generational SM fermions transform differently in the UV theories. For example, three left-handed quark doublets belong to $(\mathbf{3}\,,\mathbf{2}\,,+\frac{1}{6})_{\mathbf{F}}\subset(\mathbf{4}\,,\mathbf{4}\,,0)_{\mathbf{F}}$, $(\mathbf{3}\,,\mathbf{2}\,,+\frac{1}{6})_{\mathbf{F}}^{\prime}\subset(\mathbf{4}\,,\mathbf{6}\,,+\frac{1}{4})_{\mathbf{F}}$, and $(\mathbf{3}\,,\mathbf{2}\,,+\frac{1}{6})_{\mathbf{F}}^{\prime\prime}\subset(\mathbf{6}\,,\mathbf{4}\,,-\frac{1}{4})_{\mathbf{F}}$, when one traces back their irreps in the $\mathcal{G}_{441}$ effective theory. Likewise, three $(\mathbf{\overline{3}}\,,\mathbf{1}\,,-\frac{2}{3})_{\mathbf{F}}$’s and three $(\mathbf{1}\,,\mathbf{1}\,,+1)_{\mathbf{F}}$’s all stem from distinctive $\mathcal{G}_{441}$ irreps. Thus, the flavor non-universality under the flavor-conserving neutral currents of the $\mathcal{G}_{441}$ and $\mathcal{G}_{341}$ theories can be expected. When the theory flows to the $\mathcal{G}_{331}$ along the symmetry breaking pattern in Eq. (1), all three-generational SM fermions together with their heavy partner fermions, have already transformed the same. Therefore, the flavor universality under the flavor-conserving neutral currents of the $\mathcal{G}_{331}$ theory can be expected. These features will be explicitly demonstrated in terms of the neutral currents at each symmetry breaking stage in Sec. 3.

### v2

@@ 1 Introduction
Grand Unified Theories (GUTs), with their original formulations based on the gauge groups of $\rm SU(5)$ [1] and $\rm SO(10)$ [2], were proposed to not only unify all three fundamental symmetries into one simple Lie group of $\mathcal{G}_{U}$, but to unify all Standard Model (SM) fermions into some anomaly-free irreps of the $\mathcal{G}_{U}$ as well. The Georgi-Glashow $\rm SU(5)$ GUT [1] contains chiral fermions of $3\times\left[\mathbf{\overline{5_{F}}}\oplus\mathbf{10_{F}}\right]$, and the Fritzsch-Minkowski $\rm SO(10)$ GUT [2] contains chiral fermions of $3\times\mathbf{16_{F}}$. Meanwhile, these $\rm SU(5)$ and/or $\rm SO(10)$ GUTs, with or without their supersymmetric (SUSY) extensions, are far from satisfactory, given that three-generational SM fermion masses and the Cabibbo-Kobayashi-Maskawa (CKM) mixing pattern [3, 4] were not ultimately explained. A common feature is that the fermion contents of these minimal GUTs follow the trivially repetitive flavor structure as we have observed in the SM.
@@
In a pioneering work by Georgi [5], an extension to the gauge group of ${\rm SU}(N>5)$ was proposed to embed three-generational SM fermions non-trivially. In his original proposal, any anti-symmetric rank-$k$ fermion irrep can appear at most once with the anomaly-free condition. Physically, this leads to a non-repetitive structure for both left-handed and right-handed components of chiral fermions based on an ${\rm SU}(11)$ gauge group. Given the current LHC measurements of one single SM Higgs boson so far, which have already confirmed the hierarchical Yukawa couplings of the third-generational SM quarks/leptons and displayed the evidence for the muon [6, 7], Georgi’s original proposal is quite insightful and can be further relaxed by requiring distinctive symmetry properties of three generations in the UV theory. This relies on the concept of chiral irreducible anomaly-free fermion sets (IRAFFSs) defined in Ref. [8], which reads
Definition.
A chiral IRAFFS is a set of left-handed anti-symmetric fermions of $\sum_{\mathcal{R}}m_{\mathcal{R}}\,\mathcal{F}_{L}(\mathcal{R})$, with $m_{\mathcal{R}}$ being the multiplicities of a particular fermion representation of $\mathcal{R}$. Obviously, the anomaly-free condition reads $\sum_{\mathcal{R}}m_{\mathcal{R}}\,{\rm Anom}(\mathcal{F}_{L}(\mathcal{R}))=0$. We also require the following conditions to be satisfied for a chiral IRAFFS:
  • the greatest common divisor (GCD) of the $\{m_{\mathcal{R}}\}$ should satisfy that ${\rm GCD}\{m_{\mathcal{R}}\}=1$;
  • the fermions in a chiral IRAFFS can no longer be removed, which would otherwise bring non-vanishing gauge anomalies;
  • there should not be any singlet, self-conjugate, adjoint fermions, or vectorial fermion pairs in a chiral IRAFFS.
Obviously, one SM generational fermions formulate one chiral IRAFFS, and we found that a minimal ${\rm SU}(8)$ theory [9, 8, 10] is sufficient to have all three-generational SM quarks/leptons transform differently with two distinctive chiral IRAFFSs in the UV setup  [^1] To count the SM generations by decomposing the ${\rm SU}(8)$ fermion irreps into the ${\rm SU}(5)$ irreps [5], one finds three identical SM $\mathbf{\overline{5_{F}}}$’s and three distinctive $\mathbf{10_{F}}$’s. Hence, it is sufficient to obtain three distinctive SM generations in the UV setup of the ${\rm SU}(8)$ theory..
The ${\rm SU}(8)$ symmetry breaking pattern includes three intermediate scales above the EW scale by counting its rank. By analyzing the emergent non-anomalous global $\widetilde{{\rm U}}(1)_{T}$ symmetry in the UV, we derived the non-anomalous global $B-L$ symmetry in the SM, and one unique SM Higgs doublet from the $\mathbf{70_{H}}$ of the minimal setup was conjectured based on the requirement of the vanishing global $B-L$ charge [8]. This SM Higgs doublet was found to give natural top quark mass with the renormalizable Yukawa coupling at the tree level. Due to the distinctive symmetry properties of three-generational SM fermions in the UV, we found in Ref. [10] that all SM quark/lepton masses as well as the CKM mixing pattern can be explained with (i) both $d=5$ direct Yukawa couplings and indirect Yukawa couplings to the $d=5$ irreducible Higgs mixing operators induced by the inevitable gravitational effects, (ii) three reasonable intermediate symmetry-breaking scales, and (iii) precise identifications of the SM flavors in their UV irreps.
@@
A central issue of a unified theory is to achieve the gauge coupling unification in terms of their renormalization group equations (RGEs)  [^2] Some early studies of the RGEs based on the ${\rm SU}(5)$ and ${\rm SO}(10)$ groups include Refs. [11, 12, 13].. A strong-weak-weak (SWW) symmetry breaking pattern of
@@
were predicted with the reduced Planck scale of $M_{\rm pl}=(8\pi G_{N})^{-1/2}=2.4\times 10^{18}\,{\rm GeV}$. The maximally broken subgroup of $\mathcal{G}_{441}$ further admits the $\mathcal{G}_{\rm SM}$ as its subgroup  [^3] Other different symmetry breaking patterns of the ${\rm SU}(8)$ group were previously described in Refs. [14, 15].. These scales were obtained by the reasonable predictions of three-generational SM quark/lepton masses and the CKM mixing pattern, together with the precise identifications of the SM flavors in their UV irreps. In this paper, we perform the RG analysis of the gauge couplings in the extended strong and weak sectors based on the SWW symmetry breaking pattern in Eq. (1) and the benchmark point in Eq. (2).
The rest of the paper is organized as follows. We review the ${\rm SU}(8)$ theory in section 2, with the focus on the SWW symmetry breaking pattern. Most of the contents in this section have been obtained in a previous Ref. [10]. In section 3, we define the gauge sector of the ${\rm SU}(8)$ theory along the SWW symmetry breaking pattern defined in section 2. The gauge couplings at each symmetry breaking stage will be defined. Due to the intrinsic symmetry properties, we explicitly display the flavor non-universalities of the $\mathcal{G}_{441}$ and $\mathcal{G}_{341}$ theories as well as the flavor universality of the $\mathcal{G}_{331}$ theory through their neutral currents. In section 4, we study the gauge coupling evolutions in terms of the two-loop RGEs. The ${\rm SU}(8)$ theory with the minimal fermion/Higgs contents as well as the intermediate symmetry breaking scales in Refs. [8, 10] cannot achieve the gauge coupling unification in terms of the conventional field theoretical analyses. Instead, inclusion of additional ${\rm SU}(8)$ adjoint Higgs fields and the $d=5$ Hill-Shafi-Wetterich (HSW) operator [16, 17] are suggested to achieve the gauge coupling unification within the field theoretical framework. We summarize and make future perspective in section 5, such as the possible issues in the supersymmetric extensions. Appendix A is given to define the gauge group indices and the decomposition rules according to the SWW symmetry breaking pattern in Eq. (1). Appendix B is added to prove a unique and elegant feature of the ${\rm U}(1)$ charge quantization in the ${\rm SU}(N)$ maximally symmetry breaking pattern.
2 The ${\rm SU}(8)$ theory and its SWW symmetry breaking pattern
@@ 2.1 The setup of the ${\rm SU}(8)$ theory
The ${\rm SU}(8)$ theory contains two following chiral IRAFFSs at the GUT scale [8, 10]
@@
with undotted/dotted indices for the $\mathbf{\overline{8_{F}}}$’s in the rank-$2$ IRAFFS and the rank-$3$ IRAFFS, respectively. The Roman numbers and the Arabic numbers are used for the heavy partner fermions and the SM fermions, respectively. Three $\mathbf{10_{F}}$’s from the $\mathbf{28_{F}}$ and $\mathbf{56_{F}}$ are transforming differently, while three identical $\mathbf{\overline{5_{F}}}$’s come from nine $\mathbf{\overline{8_{F}}}$’s, as one can find in Tabs. 2, 3, and 4. These facts are sufficient to guarantee that three-generational SM quarks and leptons must transform differently in the UV theory.
The non-anomalous global symmetries observed by Dimopoulos-Raby-Susskind (DRS) [18] from chiral fermions in Eq. (3) are
@@
and we also denote the anomalous global Peccei-Quinn (PQ) symmetries [19] as
@@
Both $\widetilde{{\rm U}}(1)_{T_{2}}$ and $\widetilde{{\rm U}}(1)_{{\rm PQ}_{2}}$ can be thought as the linear combinations of two global $\widetilde{{\rm U}}(1)$’s of the $\mathbf{\overline{8_{F}}}^{\omega}$ and the $\mathbf{28_{F}}$ in the rank-2 IRAFFS, and similarly for the $\widetilde{{\rm U}}(1)_{T_{3}}$ and $\widetilde{{\rm U}}(1)_{{\rm PQ}_{3}}$ symmetries in the rank-3 IRAFFS. Since the global $B-L$ symmetry should be identical for all three generations, we further require a common non-anomalous $\widetilde{{\rm U}}(1)_{T}\equiv\widetilde{{\rm U}}(1)_{T_{2}}=\widetilde{{\rm U}}(1)_{T_{3}}$ between two IRAFFSs. Likewise, we also assume a common anomalous $\widetilde{{\rm U}}(1)_{\rm PQ}\equiv\widetilde{{\rm U}}(1)_{{\rm PQ}_{2}}=\widetilde{{\rm U}}(1)_{{\rm PQ}_{3}}$ between two IRAFFSs in Eq. (5). Accordingly, the non-anomalous global $\widetilde{{\rm U}}(1)_{T}$ charges and the anomalous global $\widetilde{{\rm U}}(1)_{\rm PQ}$ charges for fermions are assigned in Tab. 1, where the anomalous global $\widetilde{{\rm U}}(1)_{\rm PQ}$ charges are assigned such that
$p:q_{2}\neq-3:+2\,,\quad p:q_{3}\neq-3:+1\,.$ (6)
@@
We prove in App. B that the maximal symmetry breaking pattern of ${\rm SU}(8)\to\mathcal{G}_{441}$ is the unique one where an additional normalization of the ${\rm U}(1)$ charge is unnecessary. The $\mathbf{63_{H}}$ can be decomposed under the $\mathcal{G}_{441}$ as follows
@@
where the VEV is developed by the $(\mathbf{1}\,,\mathbf{1}\,,0)_{\mathbf{H}}$ component, and the $(\mathbf{\overline{4}}\,,\mathbf{4}\,,+\frac{1}{2})_{\mathbf{H}}\oplus(\mathbf{4}\,,\mathbf{\overline{4}}\,,-\frac{1}{2})_{\mathbf{H}}$ represent the Nambu-Goldstone bosons. We assign the non-anomalous global $\widetilde{{\rm U}}(1)_{T}$ charges and the anomalous global $\widetilde{{\rm U}}(1)_{\rm PQ}$ charges for Higgs fields in Tab. 1 according to the renormalizable Yukawa couplings in Eq. (7). Obviously, the only non-renormalizable term in Eq. (7) is explicitly $\widetilde{\rm SU}(5)_{\dot{\omega}}$-breaking, and also has a non-vanishing global $\widetilde{{\rm U}}(1)_{\rm PQ}$ charge according to Tab. 1 and Eq. (6). Thus, it can be explicitly broken by the gravitational effect [20, 21, 22, 23, 24].
@@ 2.2 The SWW symmetry breaking pattern of the ${\rm SU}(8)$ theory
According to Ref. [25], only the adjoint Higgs field can achieve the maximal symmetry breaking pattern of ${\rm SU}(N)\to{\rm SU}(N-k)\otimes{\rm SU}(k)\otimes{\rm U}(1)$ (with $k=[\frac{N}{2}]$) in a non-SUSY gauge theory. The SWW symmetry breaking pattern in Eq. (1) follows from Ref. [10] will be further studied in the following discussions. The GUT scale symmetry breaking will be due to the VEV of the ${\rm SU}(8)$ adjoint Higgs field of $\mathbf{63_{H}}$. In the subgroup of $\mathcal{G}_{441}$, both the strong and the weak symmetries are extended beyond the SM gauge symmetries. All intermediate symmetry breaking stages will be due to the VEVs of the Higgs fields from the gauge-invariant Yukawa couplings in Eq. (7). After the first stage of symmetry breaking pattern of $\mathcal{G}_{441}\to\mathcal{G}_{341}$ in Eq. (1), the sequential symmetry breaking stages are uniquely determined. Along this SWW symmetry breaking pattern, the non-anomalous global $\widetilde{{\rm U}}(1)_{T}$ symmetry will evolve to the global $\widetilde{{\rm U}}(1)_{B-L}$ at the EW scale according to the following sequence [8]
@@ 2.3 Decompositions of the ${\rm SU}(8)$ fermions
Table 2: The $\rm SU(8)$ fermion representation of $\mathbf{\overline{8_{F}}}^{\Omega}$ under the $\mathcal{G}_{441}\,,\mathcal{G}_{341}\,,\mathcal{G}_{331}\,,\mathcal{G}_{\rm SM}$ subgroups for the three-generational ${\rm SU}(8)$ theory, with $\Omega\equiv(\omega\,,\dot{\omega})$. Here, we denote $\underline{{\mathcal{D}_{R}^{\Omega}}^{c}={d_{R}^{\Omega}}^{c}}$ for the SM right-handed down-type quarks, and ${\mathcal{D}_{R}^{\Omega}}^{c}={\mathfrak{D}_{R}^{\Omega}}^{c}$ for the right-handed down-type heavy partner quarks. Similarly, we denote $\underline{\mathcal{L}_{L}^{\Omega}=(\ell_{L}^{\Omega}\,,-\nu_{L}^{\Omega})^{T}}$ for the left-handed SM lepton doublets, and $\mathcal{L}_{L}^{\Omega}=(\mathfrak{e}_{L}^{\Omega}\,,-\mathfrak{n}_{L}^{\Omega})^{T}$ for the left-handed heavy lepton doublets. All left-handed neutrinos of $\check{\mathcal{N}}_{L}$ are sterile neutrinos, which are $\mathcal{G}_{\rm SM}$-singlets and do not couple to the EW gauge bosons.
@@
Table 3: The $\rm SU(8)$ fermion representation of $\mathbf{28_{F}}$ under the $\mathcal{G}_{441}\,,\mathcal{G}_{341}\,,\mathcal{G}_{331}\,,\mathcal{G}_{\rm SM}$ subgroups for the three-generational ${\rm SU}(8)$ theory. All SM fermions are marked with underlines.
@@
Table 4: The $\rm SU(8)$ fermion representation of $\mathbf{56_{F}}$ under the $\mathcal{G}_{441}\,,\mathcal{G}_{341}\,,\mathcal{G}_{331}\,,\mathcal{G}_{\rm SM}$ subgroups for the three-generational ${\rm SU}(8)$ theory. All SM fermions are marked with underlines.
By following the SWW symmetry breaking pattern in Eq. (1), we tabulate the fermion representations at various stages of the ${\rm SU}(8)$ model in Tabs. 2, 3, and 4. All ${\rm U}(1)$ charges at different stages are obtained according to Eqs. (233a), (233b), (233c), and (233d). According to the rule by Georgi [5], it is straightforward to find $n_{g}=3$ in the current setup. All chiral fermions are named by their $\mathcal{G}_{\rm SM}$ irreps. For the right-handed down-type quarks of ${\mathcal{D}_{R}^{\Omega}}^{c}$, they are named as follows
@@
The flavor indices in Eqs. (2.3) and (2.3) are chosen in accordance with the previous conventions in Ref. [10]. The SM flavor names in Tabs. 2, 3, and 4 have been determined according to the consistent mass hierarchies and the CKM mixing patterns.
Through Tabs. 2, 3, and 4, it is straightforward to find that all three-generational SM fermions transform differently in the UV theories. For example, three left-handed quark doublets belong to $(\mathbf{3}\,,\mathbf{2}\,,+\frac{1}{6})_{\mathbf{F}}\subset(\mathbf{4}\,,\mathbf{4}\,,0)_{\mathbf{F}}$, $(\mathbf{3}\,,\mathbf{2}\,,+\frac{1}{6})_{\mathbf{F}}^{\prime}\subset(\mathbf{4}\,,\mathbf{6}\,,+\frac{1}{4})_{\mathbf{F}}$, and $(\mathbf{3}\,,\mathbf{2}\,,+\frac{1}{6})_{\mathbf{F}}^{\prime\prime}\subset(\mathbf{6}\,,\mathbf{4}\,,-\frac{1}{4})_{\mathbf{F}}$, when one traces back their irreps in the $\mathcal{G}_{441}$ effective theory. Likewise, three $(\mathbf{\overline{3}}\,,\mathbf{1}\,,-\frac{2}{3})_{\mathbf{F}}$’s and three $(\mathbf{1}\,,\mathbf{1}\,,+1)_{\mathbf{F}}$’s all stem from distinctive $\mathcal{G}_{441}$ irreps. Thus, the flavor non-universality under the flavor-conserving neutral currents of the $\mathcal{G}_{441}$ and $\mathcal{G}_{341}$ theories can be expected. When the theory flows to the $\mathcal{G}_{331}$ along the symmetry breaking pattern in Eq. (1), all three-generational SM fermions together with their heavy partner fermions, have already transformed the same. Therefore, the flavor universality under the flavor-conserving neutral currents of the $\mathcal{G}_{331}$ theory can be expected. These features will be explicitly demonstrated in terms of the neutral currents at each symmetry breaking stage in Sec. 3.